OpenStudy (anonymous):

Solve S = 4v2 for v.

jhonyy9 (jhonyy9):

v^2 =S/4 v_1,2 =+/- sqrt(S/4) = +/- (sqrt(S))/2

OpenStudy (anonymous):

\[v=\pm \frac{\sqrt{S}}{2} \]